(x-3)^2-6=0 ???
solve each equation by using the square root property. write all radicals in simplest form. thanks so much for your help!!!

Answer by lwsshak3(11628)

(x-3)^2-6=0 ???
solve each equation by using the square root property. write all radicals in simplest form.
(x-3)^2-6=0
(x-3)^2=6
Take sqrt of both sides
x-3=±√6
x=3+√6
or
x=3-√6